# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ 3.1 Indicator functions and lower semicontinuous functions
Let us first characterize the sup-irreducible points. Let $f$ be a lower semicontinuous function taking finite values at two distinct points $x_{0}$ and $x_{1}$. Define two functions $\chi_{0},\chi_{1}$ by setting $\chi_{0}(x_{0})=-1$, $\chi_{1}(x_{1})=-1$ and 00 elsewhere. Then $f=\sup(f+\chi_{0},f+\chi_{1})$, with the two latter functions being lower semicontinuous as a sum of such functions, so $f$ is not sup-irreducible. The same applies if $f$ takes the +$\infty$+ ∞ value at two distinct points $x_{0}$ and $x_{1}$ by setting $f_{0}=f$ everywhere except at $x_{0}$ where we choose $f_{0}(x_{0})=0$ (resp. $f_{1}(x_{1})=0$), so $f=\sup(f_{0},f_{1})$. This shows that every sup-irreducible function has to be of the form $\delta^{\bot}_{x}+\lambda$ with $\lambda\in\overline{\mathbb{R}}$. For $\delta^{\bot}_{x}$ to be lower semicontinuous, we need to have $\mathscr{X}\backslash\{x\}$ to be closed, i.e. $x$ to be isolated.
Assume there exists an anti-isomorphism $T$. Similarly to the proof of 30, applying 13, it is easy to show that $T(\delta^{\top}_{y}+\lambda)=\delta^{\bot}_{\phi(y)}+g(y,\lambda)$ with $\phi:\mathscr{Y}\to I_{\mathscr{X}}$ bijective and $g(y,\lambda)\in\mathbb{R}$. In the same fashion, we have a bijective $\psi:\mathscr{X}\to I_{\mathscr{Y}}$. Hence $\psi\circ\phi:\mathscr{Y}\to I_{\mathscr{Y}}$ is injective, and by the Schröder–Bernstein theorem, $\mathscr{Y}$ and $I_{\mathscr{Y}}$ are thus in bijection. Considering $T^{-1}$, the sets $\mathscr{X}$ and $I_{\mathscr{X}}$ are also in bijection. So we can construct with these bijections and the previous injections, two injections from $\mathscr{X}$ to $\mathscr{Y}$ and $\mathscr{Y}$ to $\mathscr{X}$. By the Schröder–Bernstein theorem, we derive a bijection between $\mathscr{X}$ and $\mathscr{Y}$.
@@
3.2 Busemann points and 1-Lipschitz and Lipschitz functions
Sánchez and Sánchez first showed in [SS11, Theorem 1] that, for Lipschitz functions on bounded metric spaces, the structure of order isomorphisms was necessarily that of lower semicontinuous functions, obtained in 30, with a bi-Lipschitz $\phi$. Extending this result to general metric spaces, [LT16, Theorem 5.5, eq.(8)] provided an “if an only if” condition on the existence of a bi-Lipschitz $\phi$. However at the moment, to the best of our knowledge there is no sufficient condition on the form of the function $g:\mathscr{X}\times\mathbb{R}\rightarrow\mathbb{R}$, nor any specific results for order isomorphisms of the space of $1$-Lipschitz functions. [^4] On Lipschitz functions, even for $\phi=\operatorname*{Id}$, the monograph [AZ90, Chapter 7] says: “Quite amazing is the fact that the operator $J$ may act in a Hölder space, although the generating function $g$ is not continuous.” Further conditions, like global Lipschitzianity of $J$, can however ensure that $J$ is affine as in [AZ90, Theorem 7.8] for $\mathscr{X}=[0,1]$. The necessity of extra assumptions on $J$ is also discussed therein if one is interested by differentiability properties.
Nevertheless a similar result to what follows was shown in [CJ17, Corollary 4.1] for 1-Lipschitz functions, for order isomorphisms that also preserve convex combinations, i.e. $J(sf+(1-s)f^{\prime})=sJf+(1-s)Jf^{\prime}$, through very different proof techniques such as bases of topologies. The line of research of Cabello et al. on weak metrics was adapted later in [DJVM20, Theorem 3.1] to Finsler manifolds. We instead do not ask that convexity is preserved but that $J$ commutes with the addition of constants.
@@ 3.2 Busemann points and 1-Lipschitz and Lipschitz functions
If $\mathscr{X}$ is a set, we say that a map $\delta:\mathscr{X}\times\mathscr{X}\to\mathbb{R}_{\geq 0}$ is a weak metric if, for all $x,y,z\in\mathscr{X}$, $\delta(x,x)=0$, $\delta(x,y)\geq 0$ and $\delta(x,z)\leq\delta(x,y)+\delta(y,z)$, and if $\delta(x,y)=\delta(y,x)=0$ implies that $x=y$. We say that a map $f:\mathscr{X}\to\mathbb{R}$ is nonexpansive with respect to $\delta$, or $1$-Lipschitz, if $f(x)\leq\delta(x,y)+f(y)$ holds for all $x,y\in\mathscr{X}$. We denote by $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$ the set of $1$-Lipschitz maps $f:\mathscr{X}\to\mathbb{R}$.
If $\delta$ is a weak metric, then the symmetrized map $\delta_{s}(x,y)=\delta(x,y)+\delta(y,x)$ defines an ordinary metric on $\mathscr{X}$. Moreover, any map $f$ that is nonexpansive with respect to $\delta$ is nonexpansive with respect to $\delta_{s}$. For $p\in(0,1]$, by subadditivity of $t\in\mathbb{R}_{+}\mapsto t^{p}$, the result below also applies to $p$-Hölder functions since $\delta^{p}$ is also a weak metric. Other interesting examples of weak metrics are those of the form “$\inf_{\gamma}\int L(\gamma)$” with nonnegative Lagrangian $L$, used in [Vil09, Chapter 7] or [AFG23, Section 7.2].
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be the set of nonexpansive maps over a set equipped with a weak metric $(\mathscr{X},\delta)$ (resp. $(\mathscr{Y},\delta^{\prime})$). Assume that $(\mathscr{X},\delta)$ satisfies either i) that the balls of $(\mathscr{X},\delta_{s})$ are compact or ii) that $\delta$ is symmetric and $(\mathscr{X},\delta_{s})$ is complete, and that the same is true for $(\mathscr{Y},\delta^{\prime})$. Then every (max,+)-isomorphism $J$ from $\mathscr{F}$ onto $\mathcal{G}$ is of the form
@@
where $g:\mathscr{X}\to\mathbb{R}$ is nonexpansive and $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ is such that $g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$ for all $x,x^{\prime}\in\mathscr{X}$. If either $\delta$ or $\delta^{\prime}$ is a metric, then $g$ is constant and $\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$.
@@
We fix a basepoint $\bar{x}\in\mathscr{X}$. We define a map $\imath$ from $\mathscr{X}$ to the set of functions from $\mathscr{X}$ to $\mathbb{R}$ by associating to any $x\in\mathscr{X}$ the function $\imath(x)\in\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$,
@@
By the triangle inequality, $-\delta(y,\bar{x})\leq i_{x}(y)\leq\delta(y,\bar{x})$, so that $i(\mathscr{X})$ can be identified to a subset of the Cartesian product $\prod_{y\in\mathscr{X}}[-\delta(y,\bar{x}),\delta(y,\bar{x})]$ which is compact in the product topology, by Tychonoff’s theorem. The horocompactification of $\mathscr{X}$ is defined as the closure $\overline{\imath(\mathscr{X})}$ of $\imath(\mathscr{X})$ in the product topology. By Ascoli’s theorem, the closure of this set is the same if we take the topology of uniform convergence on compact sets of the metric space $(\mathscr{X},\delta_{s})$. We set $\mathscr{X}(\infty)\coloneqq\overline{\imath(\mathscr{X})}\setminus\imath(\mathscr{X})$. An element of $\mathscr{X}(\infty)$ is called an horofunction.
@@
holds for all $\alpha$ and $\beta$ large enough, with $\alpha\leq\beta$. This definition is similar to the one of Rieffel [Rie02], except that Rieffel requires an almost-geodesic to be a curve, see [AGW09, Prop. 7.12] for a comparison. Note that a restricted definition of almost-geodesics, defined as sequences, was adopted in [AGW09], the more general notion in terms of nets was introduced in [Wal18]. When the balls of $(\mathscr{X},\delta_{s})$ are compact, then, $\mathscr{X}$ is a countable union of compact sets, and so, the topology of $\bar{\imath(\mathscr{X})}$ is metrizable. Then, limits in this space can be characterized in terms of sequences, rather than of nets, and so it suffices to consider sequences that are almost-geodesic.
@@
Assume either that i) all the balls of $(\mathscr{X},\delta_{s})$ are compact or ii) $\delta$ is symmetric and $(\mathscr{X},\delta)$ is complete. Then, if $h$ is a Busemann point, then, any almost-geodesic $x_{\alpha}$ converging to $h$ is such that $\delta_{s}(\bar{x},x_{\alpha})$ tends to infinity.
@@
Suppose now that $x_{\alpha}$ is an almost-geodestic and $\delta_{s}(\bar{x},x_{\alpha})$ has a bounded subnet. Then, since $(\mathscr{X},\delta_{s})$ has compact balls by assumption, we may assume $x_{\alpha}$ converges to some $x\in\mathscr{X}$ for $\delta_{s}$, and thus for $\delta$. Then, using the above observation, $\imath(x_{\alpha})$ converges to $\imath(x)\in\imath(\mathscr{X})$, contradicting that $h=\lim_{\alpha}\imath(x_{\alpha})\in\mathscr{X}(\infty)=\overline{\imath(\mathscr{X})}\setminus\imath(\mathscr{X})$.
@@
$\delta_{s}(\bar{x},x_{\alpha})=\delta(x_{\alpha},\bar{x})+\delta(\bar{x},x_{\alpha})\leq[\delta(x_{\alpha},x_{\alpha^{\prime}})+\delta(x_{\alpha^{\prime}},\bar{x})]+[\delta(\bar{x},x_{\alpha^{\prime}})-\delta(x_{\alpha},x_{\alpha^{\prime}})+\epsilon]=\delta_{s}(\bar{x},x_{\alpha^{\prime}})+\epsilon$
So $(\delta_{s}(\bar{x},x_{\alpha}))_{\alpha}$ is almost non-decreasing, and bounded by assumption, hence it converges [Wal18, Definition 2.1, Lemma 2.4]. However the almost-geodesic property also givest, since $\delta\geq 0$,
$\delta(\bar{x},x_{\alpha})\leq\delta(\bar{x},x_{\alpha^{\prime}})-\delta(x_{\alpha},x_{\alpha^{\prime}})+\epsilon\leq\delta(\bar{x},x_{\alpha^{\prime}})+\epsilon$ (30)
so the net $(\delta(\bar{x},x_{\alpha}))_{\alpha}$ is also almost non-decreasing, and bounded, since $\delta\geq 0$ implies $\delta\leq\delta_{s}$. Consequently this net converges as well. Thus for any $\alpha\leq\beta$ sufficiently large, from (30) we obtain $\delta(x_{\alpha},x_{\beta})\leq 2\epsilon$.
@@
$0\leq\delta_{s}(x_{\alpha},x_{\alpha^{\prime}})\leq\delta_{s}(x_{\alpha},x_{\beta})+\delta_{s}(x_{\beta},x_{\alpha^{\prime}})=\delta_{s}(x_{\alpha},x_{\beta})+\delta_{s}(x_{\alpha^{\prime}},x_{\beta})\leq 4\epsilon$
So $x_{\alpha}$ is Cauchy, and thus converges. We then conclude as before. ∎
@@
Assume either that i) all the balls of $(\mathscr{X},\delta_{s})$ are compact or ii) $\delta$ is symmetric and $(\mathscr{X},\delta)$ is complete. Then, if $h$ is a Busemann point, then there is no constant $a\in\mathbb{R}$ such that $h\geq a+\delta(\cdot,\bar{x})$.
@@
and so $a-1+\delta(x_{\alpha},\bar{x})+\delta(\bar{x},x_{\alpha})=a-1+\delta_{s}(\bar{x},x_{\alpha})\leq\epsilon$. However, by 37, the almost-geodesic $x_{\alpha}$ is such that $\delta(\bar{x},x_{\alpha})$ is unbounded, so we arrive at a contradiction.∎
@@
A (max,+) isomorphism $J$ of $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X})$ must preserve inf-irreducible elements. Hence, by 36, denoting by $\mathcal{B}(\mathscr{X})$ the set of Busemann points of $\mathscr{X}$, $J$ must preserve the set $\mathbb{R}+(\imath(\mathscr{X})\cup\mathcal{B}(\mathscr{X}))$. The elements of $\imath(\mathscr{X})$ all belong to the same Archimedean class, which is maximal. By 16, a (max,+) isomorphism must send a maximal Archimedean class to a maximal Archimedean class. By 38, the Archimedean class of a Busemann point is not maximal. We conclude that $J$ preserves the set $\mathbb{R}+\imath(\mathscr{X})$.
@@
$Jf(x)-Jf(x^{\prime})=g(x)-g(x^{\prime})+f(\phi(x))-f(\phi(x^{\prime}))\leq\delta(x,x^{\prime})$ (31)
In particular, setting $f=\delta^{\prime}(\cdot,\phi(x^{\prime}))$, we get that for all $x,x^{\prime}\in\mathscr{X}$
@@
which gives the isometry property. In particular $\delta_{s}^{\prime}(\phi(x),\phi(x^{\prime}))=\delta_{s}(x,x^{\prime})$
If $\delta^{\prime}$ is a metric, then (32) gives $g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))\leq\delta(x,x^{\prime})\leq\delta_{s}(x,x^{\prime})$. Hence using that $\delta^{\prime}=\delta_{s}^{\prime}$ and $\delta_{s}^{\prime}(\phi(x),\phi(x^{\prime}))=\delta_{s}(x,x^{\prime})$, we obtain $g(x)\leq g(x^{\prime})$ for all $x,x^{\prime}\in\mathscr{X}$, so $g(\cdot)\equiv\alpha$ is constant. If $\delta$ is a metric we apply the same reasoning to (33).
@@
Example 40.

### v2

@@ 3.1 Indicator functions and lower semicontinuous functions
Let us first characterize the sup-irreducible points. Let $f$ be a lower semicontinuous function taking finite values at two distinct points $x_{0}$ and $x_{1}$. Define two functions $\chi_{0},\chi_{1}$ by setting $\chi_{0}(x_{0})=-1$, $\chi_{1}(x_{1})=-1$ and 00 elsewhere. Then $f=\sup(f+\chi_{0},f+\chi_{1})$, with the two latter functions being lower semicontinuous as a sum of such functions, so $f$ is not sup-irreducible. The same applies if $f$ takes the +$\infty$+ ∞ value at two distinct points $x_{0}$ and $x_{1}$ by setting $f_{0}=f$ everywhere except at $x_{0}$ where we choose $f_{0}(x_{0})=0$ (resp. $f_{1}(x_{1})=0$), so $f=\sup(f_{0},f_{1})$. The function $f_{0}$ is l.s.c. since its epigraph is the union of that of $f$ and of $\{x_{0}\}\times[0,+\infty]$. This shows that every sup-irreducible function has to be of the form $\delta^{\bot}_{x}+\lambda$ with $\lambda\in\overline{\mathbb{R}}$. For $\delta^{\bot}_{x}$ to be lower semicontinuous, we need to have $\mathscr{X}\backslash\{x\}$ to be closed, i.e. $x$ to be isolated.
Assume there exists an anti-isomorphism $T$. Similarly to the proof of 30, applying 13, it is easy to show that $T(\delta^{\top}_{y}+\lambda)=\delta^{\bot}_{\phi(y)}+g(y,\lambda)$ with $\phi:\mathscr{Y}\to I_{\mathscr{X}}$ bijective and $g(y,\lambda)\in\mathbb{R}$. In the same fashion, we have a bijective $\psi:\mathscr{X}\to I_{\mathscr{Y}}$. Hence $\psi\circ\phi:\mathscr{Y}\to I_{\mathscr{Y}}$ is injective, and by the Cantor–Schröder–Bernstein theorem, $\mathscr{Y}$ and $I_{\mathscr{Y}}$ are thus in bijection. Considering $T^{-1}$, the sets $\mathscr{X}$ and $I_{\mathscr{X}}$ are also in bijection. So we can construct with these bijections and the previous injections, two injections from $\mathscr{X}$ to $\mathscr{Y}$ and $\mathscr{Y}$ to $\mathscr{X}$. Again by the Cantor–Schröder–Bernstein theorem, we derive a bijection between $\mathscr{X}$ and $\mathscr{Y}$.
@@
3.2 Busemann points and 1-Lipschitz functions
In [CSCS11, Theorem 1], it was first shown that, for Lipschitz functions on bounded metric spaces, the structure of order isomorphisms was necessarily that of lower semicontinuous functions, obtained in 30, with a bi-Lipschitz $\phi$. Extending this result to general metric spaces, [LT16, Theorem 5.5, eq.(8)] provided an “if an only if” condition on the existence of a bi-Lipschitz $\phi$. However at the moment, to the best of our knowledge there is no sufficient condition on the form of the function $g:\mathscr{X}\times\mathbb{R}\rightarrow\mathbb{R}$, not even lower semicontinuity, [^5] On Lipschitz functions, even for $\phi=\operatorname*{Id}$, the monograph [AZ90, Chapter 7] says: “Quite amazing is the fact that the operator $J$ may act in a Hölder space, although the generating function $g$ is not continuous.” Further conditions, like global Lipschitzianity of $J$, can however ensure that $J$ is affine as in [AZ90, Theorem 7.8] for $\mathscr{X}=[0,1]$. The necessity of extra assumptions on $J$ is also discussed therein if one is interested by differentiability properties. nor any specific results for order isomorphisms of the space of $1$-Lipschitz functions.
Nevertheless a similar result to what follows was shown in [CSJ17, Corollary 4.1] for 1-Lipschitz functions over complete metric space, for order isomorphisms that also preserve convex combinations, i.e. $J(sf+(1-s)f^{\prime})=sJf+(1-s)Jf^{\prime}$, through very different proof techniques such as bases of topologies. The line of research of Cabello et al. on weak metrics was adapted later in [DJVM20, Theorem 3.1] to Finsler manifolds. We instead do not ask that convexity is preserved but that $J$ commutes with the addition of constants.
@@ 3.2 Busemann points and 1-Lipschitz functions
If $\mathscr{X}$ is a set, we say that a map $\delta:\mathscr{X}\times\mathscr{X}\to\mathbb{R}$ is a weak metric if, for all $x,y,z\in\mathscr{X}$, $\delta(x,z)\leq\delta(x,y)+\delta(y,z)$, and if the symmetrized map $\delta_{s}(x,y)=\delta(x,y)+\delta(y,x)$ defines an ordinary metric on $\mathscr{X}$. We say that a map $f:\mathscr{X}\to\mathbb{R}$ is nonexpansive with respect to $\delta$, or $1$-Lipschitz, if $f(x)\leq\delta(x,y)+f(y)$ holds for all $x,y\in\mathscr{X}$. We denote by $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$ the set of $1$-Lipschitz maps $f:\mathscr{X}\to\mathbb{R}$.
Unlike [CSJ17, DJVM20], we do not assume the weak metric to be nonnegative-valued, which allows us to consider the Funk metric in 40, for which $\delta_{s}$ is the Hilbert metric. We still assume that $\delta_{s}$ is a metric, hence has nonnegative values.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be the set of nonexpansive maps over a set equipped with a weak metric $(\mathscr{X},\delta)$ (resp. $(\mathscr{Y},\delta^{\prime})$). Assume that $(\mathscr{X},\delta)$ satisfies either i) that the closed balls of $(\mathscr{X},\delta_{s})$ are compact or ii) that $\delta=\frac{1}{2}\delta_{s}$ and $(\mathscr{X},\delta_{s})$ is complete, and that the same is true for $(\mathscr{Y},\delta^{\prime})$. Then every (max,+)-isomorphism $J$ from $\mathscr{F}$ onto $\mathcal{G}$ is of the form
@@
where $g:\mathscr{X}\to\mathbb{R}$ is nonexpansive and $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ is such that $g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$ for all $x,x^{\prime}\in\mathscr{X}$. If both $\delta$ and $\delta^{\prime}$ are metrics, then $g$ is constant and $\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$.
For $p\in(0,1]$ and $\delta(x,y):\mathscr{X}\times\mathscr{X}\to[0,+\infty)$, by subadditivity of $t\in\mathbb{R}_{+}\mapsto t^{p}$, the result above also applies to $p$-Hölder functions since $\delta^{p}$ is also a weak metric. Other interesting examples of weak metrics are those of the form “$\inf_{\gamma}\int L(\gamma)$” with nonnegative Lagrangian $L$, used in [Vil09, Chapter 7] or [AFG24, Section 7.2].
@@
Fix a basepoint $\bar{x}\in\mathscr{X}$. We define a map $\imath:\mathscr{X}\to\mathbb{R}^{\mathscr{X}}$ by associating to any $x\in\mathscr{X}$ the function $\imath(x)\in\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$,
@@
By the triangle inequality, $-\delta(y,\bar{x})\leq\imath(x)(y)\leq\delta(y,\bar{x})$, so that $i(\mathscr{X})$ can be identified to a subset of the Cartesian product $\prod_{y\in\mathscr{X}}[-\delta(y,\bar{x}),\delta(y,\bar{x})]$ which is compact in the product topology, by Tychonoff’s theorem. The horocompactification of $\mathscr{X}$ is defined as the closure $\overline{\imath(\mathscr{X})}$ of $\imath(\mathscr{X})$ in the product topology. As a special case of a “second” Ascoli’s theorem [Sch70, p310], since $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$ with the pointwise convergence topology is isomorphic to itself with the compact-open topology and thus with uniform convergence on compact sets, the closure of $\imath(\mathscr{X})$ is the same if we take the topology of uniform convergence on compact sets of the metric space $(\mathscr{X},\delta_{s})$. We set $\mathscr{X}(\infty)\coloneqq\overline{\imath(\mathscr{X})}\setminus\imath(\mathscr{X})$. An element of $\mathscr{X}(\infty)$ is called a horofunction.
Following [Wal18, Definition 2.1], we say that a net $\{z_{\alpha}\}_{\alpha\in\mathscr{A}}\subset\mathbb{R}$ is almost non-decreasing if, for any $\epsilon>0$, there exists $A$ such that $z_{\alpha}\leq z_{\alpha^{\prime}}+\epsilon$ for all $\alpha,\alpha^{\prime}\in\mathscr{A}$ such that $A\leq\alpha\leq\alpha^{\prime}$. As easily seen, taking limits, we have $\limsup_{\alpha}z_{\alpha}\leq\liminf_{\alpha}z_{\alpha}-\epsilon$ for all $\epsilon>0$, so bounded almost non-decreasing nets converge in $\mathbb{R}$.
@@
holds for all $\alpha$ and $\beta$ large enough, with $\alpha\leq\beta$. This definition is similar to the one of Rieffel [Rie02], except that Rieffel requires an almost-geodesic to be a curve, see [AGW09, Prop. 7.12] for a comparison. Note that a restricted definition of almost-geodesics, defined as sequences, was adopted in [AGW09], the more general notion in terms of nets was introduced in [Wal18, Definition 2.2]. When the closed balls of $(\mathscr{X},\delta_{s})$ are compact, then, $\mathscr{X}$ is a countable union of compact sets, and so, the topology of $\overline{\imath(\mathscr{X})}$ is metrizable. Then, limits in this space can be characterized in terms of sequences, rather than of nets, and so it suffices to consider sequences that are almost-geodesic.
@@
Assume either that i) all the closed balls of $(\mathscr{X},\delta_{s})$ are compact or ii) $\delta=\frac{1}{2}\delta_{s}$ and $(\mathscr{X},\delta_{s})$ is complete. Then, if $h$ is a Busemann point, then, any almost-geodesic $x_{\alpha}$ converging to $h$ is such that $\delta_{s}(\bar{x},x_{\alpha})$ tends to infinity.
As already discussed in 24, inf-stable complete subspaces of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ are sets of non-expansive functions for the $\overline{\mathbb{R}}$-valued weak metric given by $e_{x}$, so 37 can be easily adapted to this abstract setting.
@@
Suppose now that $x_{\alpha}$ is an almost-geodestic and $\delta_{s}(\bar{x},x_{\alpha})$ has a bounded subnet. Then, since $(\mathscr{X},\delta_{s})$ has compact balls by assumption, we may assume $x_{\alpha}$ converges to some $x\in\mathscr{X}$ for $\delta_{s}$. Then, using the above observation, $\imath(x_{\alpha})$ converges to $\imath(x)\in\imath(\mathscr{X})$, contradicting that $h=\lim_{\alpha}\imath(x_{\alpha})\in\mathscr{X}(\infty)=\overline{\imath(\mathscr{X})}\setminus\imath(\mathscr{X})$.
@@
$\delta_{s}(\bar{x},x_{\alpha})=\delta(x_{\alpha},\bar{x})+\delta(\bar{x},x_{\alpha})\leq[\delta(x_{\alpha},x_{\alpha^{\prime}})+\delta(x_{\alpha^{\prime}},\bar{x})]+[\delta(\bar{x},x_{\alpha^{\prime}})-\delta(x_{\alpha},x_{\alpha^{\prime}})+\epsilon]=\delta_{s}(\bar{x},x_{\alpha^{\prime}})+\epsilon.$
So $(\delta_{s}(\bar{x},x_{\alpha}))_{\alpha}$ is almost non-decreasing, and bounded by assumption, hence it converges.
@@
$0\leq\delta_{s}(x_{\alpha},x_{\alpha^{\prime}})\leq\delta_{s}(x_{\alpha},x_{\beta})+\delta_{s}(x_{\beta},x_{\alpha^{\prime}})\leq 4\epsilon.$
So $x_{\alpha}$ is Cauchy, and thus converges. We then conclude as in Case (i). ∎
@@
Assume either that i) all the closed balls of $(\mathscr{X},\delta_{s})$ are compact or ii) $\delta=\frac{1}{2}\delta_{s}$ and $(\mathscr{X},\delta_{s})$ is complete. Then, if $h$ is a Busemann point, then there is no constant $a\in\mathbb{R}$ such that $h\geq a+\delta(\cdot,\bar{x})$.
@@
and so $a-1+\delta(x_{\alpha},\bar{x})+\delta(\bar{x},x_{\alpha})=a-1+\delta_{s}(\bar{x},x_{\alpha})\leq\epsilon$. However, by 37, the almost-geodesic $x_{\alpha}$ is such that $\delta_{s}(\bar{x},x_{\alpha})$ is unbounded, so we arrive at a contradiction.∎
@@
A (max,+) isomorphism $J$ of $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X})$ must preserve inf-irreducible elements. Hence, by 36, denoting by $\mathcal{B}(\mathscr{X})$ the set of Busemann points of $\mathscr{X}$, $J$ must preserve the set $\mathbb{R}+(\imath(\mathscr{X})\cup\mathcal{B}(\mathscr{X}))$. Owing to the triangular inequality, the elements of $\imath(\mathscr{X})$ all belong to the same Archimedean class, which is maximal. By 16, a (max,+) isomorphism must send a maximal Archimedean class to a maximal Archimedean class. By 38, the Archimedean class of a Busemann point is not maximal. We conclude that $J$ preserves the set $\mathbb{R}+\imath(\mathscr{X})$.
@@
$Jf(x)-Jf(x^{\prime})=g(x)-g(x^{\prime})+f(\phi(x))-f(\phi(x^{\prime}))\leq\delta(x,x^{\prime}).$ (32)
Setting $f=\delta^{\prime}(\cdot,\phi(x^{\prime}))$, since $02\delta^{\prime}(\phi(x),\phi(x))=\delta^{\prime}(\phi(x),\phi(x))=$ 0, we get that for all $x,x^{\prime}\in\mathscr{X}$
@@
which gives the almost isometry property. Commuting the role of $x$ and $x^{\prime}$ in (35) and summing, we derive that $\delta_{s}^{\prime}(\phi(x),\phi(x^{\prime}))=\delta_{s}(x,x^{\prime})$. If $\delta$ and $\delta^{\prime}$ are metrics, then $g(x)-g(x^{\prime})=0$, so $g(\cdot)\equiv\alpha$ is constant.
@@
Example 40 (Isomorphisms of Hilbert and Funk distances).